OpenStudy (anonymous):

1.)Choose the equation of the line passing through the point (-4, -2) and perpendicular to y = -x + 6. y = -x + 6 y = x - 6 y = -x - 2 y = x + 2 2.)A video game club charges a fixed annual membership fee of $20 and $2 per video game rented. Let f(n) represent the total annual cost of renting n video games. Which of the following functions best represents the relationship between f(n) and n if the membership was increased by $15 the next year? a. f(n) = 2n + 35 b. f(n) = 20n + 2 c. f(n) = 10n + 35 d. f(n) = 2n + 10

OpenStudy (shamil98):

perpendicular lines have inverse slopes lets say the slope of a line is 3 then its perpendicular line's slope would be -1/3

OpenStudy (anonymous):

Explain more please? :O

OpenStudy (shamil98):

y = - x + 6 the coordinates for the perpendicular line is (-4,-2) the inverse of -1 is 1 so you plug it into point-slope form y + 2 = 1 (x+4) y + 2 = x + 4 y = x +2

OpenStudy (shamil98):

point-slope = y- y1 = m(x-x1)

OpenStudy (shamil98):

after point - slope i changed the equation to slope-intercept.

OpenStudy (shamil98):

y = mx + b is slope intercept

OpenStudy (anonymous):

*Do i have to distribute?

OpenStudy (shamil98):

Distribute 1? anything times 1 is that number.

OpenStudy (anonymous):

I'm confused :/ what numbers am i plugging in for the equation y- y1 = m(x-x1)

OpenStudy (shamil98):

the coordinates given (-4,-2)